Mrs. DiNardo's physics class has 115 students, classified by academic year and gender, as illustrated in
the following table. Mrs. DiNardo randomly chooses one student to collect yesterday's work. What is the
probability that she selects a female, given that she chooses randomly from only the juniors? Express
your answer as a fraction in lowest terms or a decimal rounded to the nearest millionth.
Mrs. DiNardo's Physics Class
Classification Males Females
Freshmen 20 Males 9 Females
Sophomores 18 Males 15 Females
Juniors 10 Males 12 Females
Seniors 12 Males 19 Females

Found 2 solutions by greenestamps, stanbon:
Answer by greenestamps(13200)



"...given that she chooses randomly from only the juniors..." means she is only choosing from among the 22 juniors. 12 of the 22 juniors are female, so the probability is 12/22 = 6/11.

Formally (the way conditional probability is usually taught),
                        P(junior AND female)   12/115   12    6
     P(female|junior) = -------------------- = ------ = -- = --
                            P(junior)          22/115   22   11


Answer by stanbon(75887)

Mrs. DiNardo's physics class has 115 students, classified by academic year and gender, as illustrated in
the following table. Mrs. DiNardo randomly chooses one student to collect yesterday's work. What is the probability that she selects a female, given that she chooses randomly from only the juniors?
Express your answer as a fraction in lowest terms or a decimal rounded to the nearest millionth.
Mrs. DiNardo's Physics Class
Classification Males Females
Freshmen 20 Males 9 Females
Sophomores 18 Males 15 Females
Juniors 10 Males 12 Females
Seniors 12 Males 19 Females
----
Ans: 12/22 = 6/11
---------
Cheers,
Stan H.
----------